Hi please help me and explain thorouly on how to find the equation of the line that contains: Point (0,5), M=2.(PLEASE SHOW ME STEP BY STEP CAUSE IM STUCK ON THIS) And one more problem; Is the line x= -2 perpendicular to the line y= 3? Please expalin!
Thanks I really appreciate it.

Answer by ankor@dixie-net.com(22740)

find the equation of the line that contains: Point (0,5), M=2.
:
When you are given the point and the slope use the "point/slope formula", which is:
y - y1 = m(x - x1)
:
in your problem: x1=0, y1=5, m=2, substituting in the above equation:
y - 5 = 2(x - 0)
y - 5 = 2x - 0; multiply what's inside the brackets by 2
or just
y - 5 = 2x
add 5 to both sides
y = 2x + 5; the equation of the given line
:
And one more problem; Is the line x= -2 perpendicular to the line y= 3? Please explain
when x is given a single value, it is a vertical line going thru -2
when y is given a single value, it is a horizontal line going thru +3
A vertical line is perpendicular to a horizontal line, right??
:
Did this shed some light on this stuff for you?